Tuesday, 21 December 2010


We have just been informed of a BT outage affecting 20CN connections.

The exchange currently having issues is "20CN-BRAS-ERX15-READING3".

BT are currently investigating this outage.


This issue has been cleared and all circuits are now back online.


We have received this update/cause from BT:

IMT35773/10 (F) - 21/12/2010 09:26 - 21/12/2010 12:00 - : READING 3 EDGE ROUTE SWITCH 15 : LOSS OF SERVICE : Loss of internet services for 20708 broadband end users. Cause corrupt data - service restored"

Apologies for any inconvenience caused

Monday, 15 November 2010

[CLOSED] Delayed BT updates

15/11/10 - Delayed BT updates

Due to BT system issues we are currently unable to get BT fault updates.

Apologies for any inconvenience caused, all updates will be forwarded to all our customer as soon as possible


16/11/10 - We are still experiencing intermittent issues with BT`s fault systems so some updates may be a bit delayed.

Apologies for the inconvenience, updates will be forwarded as soon as possible


BT have now repaired their fault report system so updates are back to normal

Wednesday, 3 November 2010

[Closed] - EDINBURGH Lines down

Alot of lines that connect to Edinburgh RASs went down at 11am
This does not seem to be isolated to a specific exchange or RAS.

BT have been informed and are currently investigating.

Next update expected 12:45


12:40: BT Engineers are currently onsite at the affected BRAS working on the issue


13:30: Still no official line from BT but we are seeing circuit reconnect slowly


14:20: BT are hoping to have this resolved shortly
This issue is mentioned here: http://www.theregister.co.uk/2010/11/03/bt_outage/


17:30: Unfortunatley no further updates from BT but we are seeing the majority of lines reconnected and online.

Customers that are having difficulty accessing broadband are advised to reboot their equipment.
We will post further updates when recieved


We have been informed this issue has now been resolved, once again this was due to BT hardware failure.

We also have the following form BT:

"BT’s engineers, working with specialists from BT’s network equipment vendor,  are pro-actively monitoring both the network and systems stability of the affected node. In parallel, plans are being developed to implement total resilience to this environment".

Friday, 29 October 2010

[Closed] - 20cn ADSL Connectivity

Dear Customer,

We are currently investigating issues with 20cn ADSL connectivity,

Some users may have lost connection.

Once we have an update from our engineers an update will be sent.


The current 20cn outage has been identified as a BT nationwide outage.

We are currently awaiting details/ETA from BT Wholesale.


BT Wholesale confirm just some of the areas affected are:

    * Northern Ireland
    * Scotland
    * London
    * Newcastle

BT are unable to confirm when the next update will be at this time, but all their efforts are being put into resolving this MSO.


Engineers are on-site and working to resolve the incident. Next update is due after 21.45


Engineer is still onsite working to resolve the outage, next update is due at 22.30.


This outage has now been resolved by BT Wholesale, and we can see most affected users are back online from approximately 00.30.

If you are unable to connect now please power off your router for 20 minutes, and then try to connect again once it has been powered back up.

Thank you.

Thursday, 28 October 2010

[CLOSED] - All BM-TH RASs Down

At approximatley 09:10am this morning all BM-TH RASs went down

BT are aware of the situation .

Further updates to follow shortly


10:30am - 21CN lines (21CN-BRAS-REDx--BM-TH RASs) are still down.

Today 09:58:08 BTO have now raised an incident.

This isn't just affecting people in Birmingham

As yet we've no ETA from BT.
We'll update as soon as we have further information.



BT Engineers are still working on this issue, the next update from BT Wholesale is due at 11:40am


A BT Wholesale engineer is on-site, and has found a faulty card which is responsible for this outage. He has reset the device and is monitoring it to check services come up. Should this not restore service he will replace the card. We’ll post a further update when we get one. 


BT card issue now resolved - clear time 13:30.

Apologies for any inconvenience caused during this outage.

Tuesday, 28 September 2010

[Closed] - Caching DNS server issues

We currently have an issue with one of our caching DNS servers, this looks to be a server failure, our engineers are currently investigating this issue.

Customers may see some slow responses whilst resolving DNS.

Updates will follow.


A power cycle of this server resolved the issue, there were no reported incidents of slowdown whilst using the alternative servers.

Friday, 24 September 2010

BT Line card failure (IVYBRIDGE & KINGSKERSWELL exchanges)

There is a MSIP Line card failure that is being investigated by BT at the moment. This has been classed as a major outage and BT are still investigating. We have not been given the expected restore time yet.

Affected phone codes: 01803 and 01752

Thursday, 23 September 2010

[CLOSED] - Network Outage

14:50 - We have just experienced a short network outage, this lasted approximately 1 minute, the cause of the outage is currently under investigation.


15:05  - Engineers confirm the issue has cleared. Initial investigations indicate an issue on one of our transit providers.

More information will follow shortly


17:00 - Issue is now cleared and confirmed as a Transit provider issue.
Apologies for any inconvenience caused during this period.
Delay was due to the time for BGP to reload its routes


Tuesday, 14 September 2010

[CLOSED] - Phone System Issues


Dear Customer

We are currently experiencing issues with our phone system recieving calls.

Please accept our sincere apologies for any inconvenience caused.

Any urgent issues please email support@186k.co.uk


13:00 Phone system issues are now fixed by our engineers.
Again, please accept our sincere apologies for any inconvenience caused


Friday, 3 September 2010


21CN WBC : MILTON KEYNES BRAS RED 5 : LOSS OF SERVICE:Service affecting:Total Loss

Incident Ref: IMT24756/10

Target Completion Timestamp: 14:00 03/09/2010

Further updates when recieved


2pm: Issue cleared and customers now online
Apologies for any inconvenience


Tuesday, 17 August 2010


Circuits connecting to BT 21CN-BRAS-RED3-SL have been having connectivity issues

Incident Ref: IMT23225/10

Incident Description: 21CN WBMC : SLOUGH : LOSS OF SERVICE:Service affecting:Total Loss

Target Completion: 17-08-2010 20:30:00


All affected circuits are now online


Wednesday, 11 August 2010



Start Date: 11/08/2010 13:35:00 Estimated Completion Date: 11/08/2010 17:00:00.

Incident Reference: IMT22571/10
21CN WBMC : SOUTHBANK BRAS RED7 : LOSS OF SERVICE:Service affecting:Total Loss


17:30 - Customers are now online. Apologies for any inconvenience caused


Tuesday, 10 August 2010

[CLOSED] - Network Outage

Between 2:30am and 6am we experienced some network issues and some customer will have experienced intermittent service.

It was triggered by work done by NTL and caused an issue with one of our switches in Telehouse.

Service is now fully restored, apologies for any inconvenience caused

The 186k Team

Tuesday, 27 July 2010

CLOSED - Phone System Issue

Dear customer,

We are currently experiencing issues with our phone system, this is under investigation.

We are still contactable via email (support@186k.co.uk)

Please accept our apologies for any inconvenience caused.


Dear customer,

Our phone system is now fully operational again.

Please accept our apologies for any inconvenience caused during this short outage.

Thursday, 22 July 2010

[CLOSED] - 21CN BRAS RED1 - North/West Yorkshire

We are seeing multiple lines down in and around the Leeds area (including Bradford, Dewsbury,Cleckheaton, Harrogate)

We have reported this to BT, who are currently investigating.

Apologies for the inconvenience, further updates will be posted ASAP


14:00 BT are still working on this issue and have assured us they will have an update around 15:00


15:00 We are now seeing all the connections online related to this outage.
More information will be provided when recieved from BT


Friday, 9 July 2010

[CLOSED] - Slow response/ Can't connect for Business users in Northern Ireland and Scotland

BT are aware of a temporary problem in the Scotland and Northen Ireland areas that may be causing some users difficulties accessing the Internet. bt engineers are working to resolve this problem and we apologise for any inconvenience this may cause.

Further updates when recieved from BT


13:15 - BT are still working to resolve this major issue
BT have confirmed that some lines on 21CN-RED5-GW-TH, 21CN-RED4-GW-TH, 21CN-RED3-GW-TH, 21CN-RED1-GW-TH are also affected.


16:00 - BT have now cleared this fault. Apologies for any inconvenience caused


Tuesday, 6 July 2010


At 13:20 all lines on the 21CN-BRAS-RED2-GI-B RAS went down.

BT are aware of this issue and are currently investigating
Further updates will follow ASAP


16:00 - Still no update from BT but we are assured they are
working to resolve the issue ASAP


12:00 - 07 July 2010

All reported lines are now connected and online.
BT have not yet advised the cause of this issue.


Wednesday, 30 June 2010


BT Wholesale are currently experiencing a major national fault meaning that users are unable to connect, or once connected may disconnect. They are working on this matter as a priority and will resolve it as quickly as possible.

We are expecting a BT Wholesale update at 10.30am.
We apologise to any customer affected by this.

Update from BT Wholesale:

This issue is still being worked on by BT engineers, the current fix ETA. is 13.00.

We will notify you of any further updates.

BT work continues, although most customer services have been restored.

We are awaiting a further update from BT at approx 15.00


BT Wholesale have confirmed this fault has now been resolved, and on their behalf we applogise to everyone who was affected by this.

Replacement hardware is on site and will be installed during an emergency maintenance window which is scheduled to take place between 00:00 and 06:00 tomorrow morning. Affected customers are expected to experience a disruption of service whilst this work is being carried out.


Apologies, we are seeing a reoccurence of the issue.
The maintenance work between 00:00 and 06:00 1st July should resolve this issue


Last nights work has resolved this issue.
Again we apologise for any inconvenience caused by this BT outage


Thursday, 10 June 2010

[CLOSED] - LLU Broadband Connectivity Issue

We are currently investigating issues with LLU connectivity.

This will only affect users on the LLU network.

We have just received the following report from our LLU provider:

"At approximately 10:30 (BST) today our LAC at THN stopped working and dropped all traffic routing through it. We are unable to contact it on its management interface and believe it may have crashed.

All traffic failed over to the LACs at TCH. The LACs at TCH are running at less than half load each, and are more than capable of handling all the network traffic on their own. However, we are now operating in a failover scenario so the network is at risk from a further failure.

We have despatched an engineer to THN to debug this locally and will update again later"

Once we receive the further update we will update this post.


Here is the update from our supplier:

"Our engineer on site was able to console into the LAC and establish that it had not crashed, but had lost arp layer communication.

We carried out various troubleshooting, including replacing SFPs and moving patches over to different router ports, but were unable to satisfactorily clear the fault or establish any cause from debugs. We have given the LAC a cold reboot and it is now performing normally.

We have brought the LAC back onto the network and allowed it to start taking some load again. We will be monitoring this device very closely now for any signs of an underlying software or hardware fault"

Please accept our apologies for any inconvenience caused.

Tuesday, 1 June 2010

[Closed] - Email Delays

Some customers are reporting delays in sending and receiving emails.

We are currently investigating this issue and will post our findings here.

Wednesday, 7 April 2010

Closed - ADSL Network Outage

We are currently experiencing intermittent issues with ADSL connectivity.

Users may be seeing the BT RAS holding page, anyone experiencing this issue please power down your router/modem for 20 - 30 minutes.

Once we have an update from BT we will update this post.

Support Team.


Update (11:00am): All DSL connections should be operational again (please reboot equipment to re authenticate if required), the issues was between the BT RAS and our Radius servers and is still under investigation, once we have the root cause we will post the update.

Hosting Services Outage.

We have experienced a network issue this evening which may have resulted in a degraded service for some users, a switch within the core network has been identified as the cause and has now been isolated.

Some users may experience slight delays on messages sent during the outage, normal service has now resumed please accept our apologies for any inconvenience this has caused.

Wednesday, 31 March 2010

Closed - 21cn BRAS issue bras-red2.bm-th

All connections affected by the outage were back online at approximately 11:15am, however we left the fault open to monitor.

We are still awaiting the official BT report.

Please accept our apologies for any inconvenience caused.


10:52 - We can see circuits are now starting to connect, could we please ask any customers to reboot their modem/router.

We are still awaiting a report from BT.


There is currently an issue with the BRAS-RED2.BM-TH (Birmingham Area).

This was due to BT planned engineering works, there were issues with the works and BT are currently investigating.

We will post any further news when updates are received.

Tuesday, 23 March 2010

Closed - DSL Radius Issues

Some customers may be experiencing DSL radius rejections, we are currently looking into this.

Once we have an update we will update this post.


This issue has been resolved.

Core Network Maintenance - 23th/24th March 2010

This evening we will be carrying out some work on the core network, we are not expecting any issues.

This work is due to commence at 00:00 and be completed at approximately 01:00

Affected service risk is: Low

Monday, 8 March 2010

Closed - ADSL DNS Outage

Some ADSL customers may have experienced a DNS outage between 3 am - 4 am this morning.

This was due to a power issue in Telehouse (currently under investigation).

If any customers are still experiencing DNS issues on their ADSL connection could you please reboot your equipment.

Thank you.

Monday, 1 March 2010

Stalled - Telinco Covs Platform

There are currently issues with the COVS hosting platform, we are working with the Telinco engineers to help them resolve this issues.

Customers on the platform will experience web and mail issues.

This issue in not affecting the 186k hosting platform.


We have moved most customers away from the Telinco platform to ensure a more reliable hosting solution.

Any customers still affected please contact us on:

Phone: 08701 222 186
Email: support@186k.co.uk

Thank you.

Thursday, 18 February 2010

Information - Transit provider issue.

We have identified an issue with one of our transit providers and this is being investigated.

We are not expecting customers to see any effects of this.

Tuesday, 9 February 2010

Closed - ADSL Intermittent connectivity

Some customers are experiencing intermittent connectivity.

This is currently under investigation with our engineers.


All affected connections are now believed to be stable.

We are still awaiting an update as to what the issue was and will add this to the post when received.

Please accept our apologies for any inconveniences caused during the issue.

Tuesday, 2 February 2010

Closed - Shared FormMail script

There is currently an issue with the shared frommail script.

Update:- 3rd Feb 10

We believe this issue will be resolved before end of play today.

FormMail issues now resolved.

Thursday, 28 January 2010

Closed - Website Hosting Issues

29th Jan - 16:00

All websites have now been restored and should now be live, if anyone is experiencing issues please email support@186k.co.uk and we will look into this.

Once again please accept our apologies for any inconvenience caused.

29th Jan - 12:26pm

A majority of the effected sites are now back online, we have an ETA for 16:00 for all sites to be back online.

29th Jan

It has been discovered that the current filer has a major fault, the standby filer is being prepared to be brought online.

Once an ETA has been confirmed I will update this post.


Unfortunately we are experiencing issues very similar to this morning, our engineers are currently working on this issue.

Updates will follow once received.


The issue was due to to connections to the filer, to resolve this issue the filer required a reboot, this has now been complete and all websites will now be operational.

Please accept our apologies for any inconvenience caused during this issue.


We are currently experiencing issues with some shared hosted web sites.

This is being investigated by our engineers and once we receive an update we will update this notice.

Monday, 25 January 2010

Closed - Monitoring - Pop3/SMTP Email Issue

26th Jan 2009 - 10:21

Yesterday we removed a node (Node-2) from the cluster due to sporadic issues during the collection of email (POP3). This node was repaired and tested before being reintroduced, however the same node has developed issues with the SMTP service, this node has been completely removed from the cluster to undergo repairs. Please accept our apologies for any inconvenience caused during this issue.


This morning some customers experienced issues whilst trying to collect email, sometimes the username/password was rejected, this was due to an issue with one of the email nodes (connecting to any other node would have resulted in a successful pop of the mailbox).

To rectify this issue we have removed the node that was not allowing connections.

Please accept our apologies for any inconvenience caused.


There have been further reports of intermittent Mail issues, this is currently being investigated.

Tuesday, 19 January 2010

Closed - Automated ADSL Diagnostics

We are currently experiencing issues with the Dolphin portal Automated ADSL Diagnostics,the results that are being returned are incorrect. This is currently being investigated by our engineers and this post will be updated once the a resolve has been implemented. Please accept our apologies for any inconvenience caused during this issue.

Thursday, 14 January 2010

Closed - 21cn - bras-red2.sl

We have seen some customer intermittent connections in the Slough area (Bras-red2.sl).

We are currently speaking to BT regarding this issue, and will update this page once we have some more information.

Monday, 4 January 2010

Closed - Website / Email Issues

Some customers are experiencing slow email (sending/receiving) and slow webpages.

Our engineers are currently investigating this issue.

Once we know the cause of this issue we will update the status page.